HIGH RESOLUTION SPECTROSCOPY OF B EXCITED STATE OF BENZENE DIMER CATION

K. FURUKAWA, T. KAWASAKI, T. KITAGAWA, E. SHIROTA, T. ICHIYANAGI, S. NISHI, O.A. MATSUZAKI, and \underline{H. SATO}, Department of Chemistry for Materials, Faculty of Engineering, Mi’e University, Tsu 514-8507, Japan.

High resolution spectrum of B-X absorption band of benzene dimer cation in the 600-700 nm region ^{a} was measured with a Coherent 899-21/899-29 ring dye laser, by monitoring predissociation in an Extrel triple quadrupole mass spectrometer. The spectrum was featured with a ca. 200-cm^{−1} progression.

aK. Ohashi and N. Nishi, J. Chem. Phys.95, 4002 (1991); K. Ohashi, Y. Nakai, T. Shibata, and N. Nishi, Laser Chem.14,3 (1994)